# GLIP-OOD: Zero-Shot Graph OOD Detection with Foundation Model

### v1

@@ Abstract
Out-of-distribution (OOD) detection is critical for ensuring the safety and reliability of machine learning systems, particularly in dynamic and open-world environments. In the vision and text domains, zero-shot OOD detection—which requires no training on in-distribution (ID) data—has made significant progress through the use of large-scale pretrained models such as vision-language models (VLMs) and large language models (LLMs). However, zero-shot OOD detection in graph-structured data remains largely unexplored, primarily due to the challenges posed by complex relational structures and the absence of powerful, large-scale pretrained models for graphs. In this work, we take the first step toward enabling zero-shot graph OOD detection by leveraging a graph foundation model (GFM). We show that, when provided only with class label names, the GFM can perform OOD detection without any node-level supervision—outperforming existing supervised methods across multiple datasets. To address the more practical setting where OOD label names are unavailable, we introduce GLIP-OOD, a novel framework that employs LLMs to generate semantically informative pseudo-OOD labels from unlabeled data. These labels enable the GFM to capture nuanced semantic boundaries between ID and OOD classes and perform fine-grained OOD detection—without requiring any labeled nodes. Our approach is the first to enable node-level graph OOD detection in a fully zero-shot setting, and achieves state-of-the-art performance on four benchmark text-attributed graph datasets.
@@ 1 Introduction
Out-of-distribution (OOD) detection [31, 14, 9] is the task of identifying inputs that deviate significantly from the training distribution. As machine learning models are increasingly deployed in dynamic and safety-critical environments, the ability to reliably detect OOD instances has become important for maintaining robustness, improving trustworthiness, and preventing catastrophic failures. Applications span many domains, including medical diagnosis [23], autonomous driving [6], video analytics [11], and financial anomaly detection [2], where encountering unseen or unexpected data is common and consequential.
@@
Graph OOD Detection and Current Landscape. Recently, OOD detection has gained traction in graph-structured data (known as graph OOD detection [27, 15, 29, 22]), where the goal is to detect anomalous or previously unseen nodes, edges, or subgraphs. It is important to the domains such as social networks [28], citation graphs, and e-commerce platforms, where new entities (e.g., emerging users, novel research topics, or innovative products) may appear during inference. These graph OOD detection methods generally require a substantial number of labeled in-distribution (ID) nodes to train a graph neural network (GNN)-based classifier, which subsequently produces class probability distributions (logits) for all nodes in the graph [27, 15, 29]. OOD detection is then performed in a post-hoc manner by applying scoring functions—such as energy-based [27], maximum softmax probability [9], or uncertainty-based metrics [15]—to these logits. Nodes with high energy or entropy scores are considered likely to be OOD samples. However, these methods are fundamentally limited by their reliance on labeled ID data, which can be expensive and impractical to obtain in real-world graph scenarios. This raises a natural question: Can we perform graph OOD detection without relying on labeled nodes at all?
Figure 1: GLIP-OOD-R denotes using real labels, while GLIP-OOD-L uses LLM-generated labels. When provided with only label names and no node-level supervision, GLIP-OOD achieves OOD detection performance that surpasses all supervised graph learning methods trained on large amounts of ID-labeled data across all datasets.
@@
Zero-Shot OOD Detection in Vision and Text. In computer vision and NLP areas, zero-shot OOD detection—which requires no training on ID data—has emerged as a compelling research direction. In the image domain, methods [24, 18, 5, 1] leverage the capabilities of large-scale vision-language models (VLMs), e.g., CLIP [19], to detect OOD samples across a variety of ID datasets without requiring any labeled training examples. For example, CLIPN [24] enhances CLIP with the ability to distinguish between OOD and ID samples by leveraging both positive semantic prompts and negation-based prompts, enabling zero-shot OOD detection in the image domain. ZOC [5] extends CLIP by dynamically generating candidate unseen labels per test image and introducing a confidence score based on feature-space similarity to both seen and generated unseen labels. In the text domain, large language models (LLMs) have demonstrated impressive zero-shot capabilities by leveraging natural language prompts and contextual understanding to detect OOD content [32, 10]. The ability to perform effectively with limited data makes LLM-based OOD detection particularly attractive for real-world applications. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored due to the unique challenges posed by relational structures. More importantly, until now, there has been no powerful large-scale foundation model—like VLMs and LLMs—for the graph domain.
Our Proposal: Zero-Shot OOD Detection in Graphs. In this paper, we take the first step toward enabling zero-shot graph OOD detection using a graph foundation model (GFM) [33, 12, 13, 3]. Similar to CLIP, the original GFM lacks built-in OOD detection capabilities—it is designed to match any given node to one of the provided seen class labels. To enable effective OOD detection, we first introduce only the label names to the GFM and directly derive the output logits from it to perform zero-shot OOD detection. As shown in Fig. 1, we empirically find that when provided solely with label names, the GFM achieves OOD detection performance that surpasses all supervised graph learning methods relying on large numbers of ID labels to train an ID classifier across all datasets. This demonstrates that the GFM is inherently capable of performing zero-shot OOD detection and holds significant potential to become highly OOD-aware, provided the ID and OOD classes are clearly defined. However, in practical open-world scenarios, we are often unable to access the actual OOD label space. To address this, we propose generating a set of pseudo-OOD labels based on the inherent knowledge within the graph and performing OOD detection on an augmented label space that includes both the original ID classes and the newly synthesized OOD labels.
In contrast to image and text OOD detection, node-level graph OOD detection typically operates in a transductive setting, where the model has access to the complete set of nodes during training, but only the labels for the ID classes are provided. Within this transductive setting, we propose leveraging LLMs to generate pseudo-OOD label names based on the initially unlabeled graph. Specifically, we begin by randomly sampling a small subset of unlabeled nodes and use the LLM to perform coarse-grained zero-shot OOD detection. The LLM is given the names of the ID classes and asked to determine whether a query node, described by its textual attributes, belongs to any of the known ID categories. If the LLM predicts that the node does not match any ID class, we consider it a pseudo-OOD node. For each such node, we further prompt the LLM to summarize its textual description and generate a representative OOD category name, thereby creating a set of pseudo-OOD labels. We then augment the original ID label set with these pseudo-OOD labels and perform zero-shot OOD detection using the GFM across the expanded label space. This approach enables the GFM to assign each node to either one of the known ID classes or a synthesized OOD category, thereby facilitating fine-grained OOD detection without requiring any real OOD labels or manual annotations.
@@
  • To the best of our knowledge, we are the first to study the problem of zero-shot graph OOD detection and to provide a comprehensive setting and framework for enabling node-level OOD detection without any node-level supervision.
  • We propose GLIP-OOD, which uses an LLM to generate pseudo-OOD labels that are semantically distant from the given ID labels but close to the actual unknown OOD labels in the transductive graph setting. We then perform zero-shot OOD detection directly with the GFM in the augmented label space.
  • Experimental results demonstrate that our proposed zero-shot graph OOD detection method outperforms all existing supervised approaches on four benchmark TAG datasets when provided with all label names but without any node-level supervision. Furthermore, when only ID label names are provided, GLIP-OOD achieves OOD detection performance comparable to that of supervised methods, which require training an ID classifier using a large number of labeled ID nodes.
@@ 2 Preliminary
Our study focus on OOD detection on TAGs. A TAG is represented as $G_{T}=(\mathcal{V},\mathbf{A},\mathbf{T},\mathbf{X})$. The set of nodes is $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, where each node is associated with raw text attributes $\mathbf{T}=\{t_{1},t_{2},\dots,t_{n}\}$. These text attributes are converted into embeddings $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$ using a pretrained model [21]. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes graph connectivity, where $\mathbf{A}[i,j]=1$ indicates an edge between nodes $v_{i}$ and $v_{j}$.
Node-level Graph OOD Detection. We have a labeled node set $\mathcal{V}_{I}$ and a large unlabeled node set $\mathcal{V}_{U}$. Each labeled node belongs to one of $K_{\text{ID}}$ known classes $\mathcal{Y}_{\text{ID}}=\{y_{k}\}_{k=1}^{K_{\text{ID}}}$, while an unlabeled node may belong to an unknown class not included in $\mathcal{Y}_{\text{ID}}$. The problem is framed in a semi-supervised, transductive setting, where we can access the full set of nodes during training but only a subset of class labels (ID classes). Our goal is, for each node $v_{i}\in\mathcal{V}_{U}$, determine whether it is from the $K_{\text{ID}}$ ID classes or from the OOD unknown classes.
Zero-shot Graph OOD Detection with All Labels. We provide the model only with the class names and do not assume access to any node labels in the graph. That is, all nodes in the graph are unlabeled, and the model directly outputs an OOD score for each input node based solely on the label names, node attributes $X$, and the adjacency matrix $A$.
Zero-shot Graph OOD Detection. Only the names of ID classes, node attributes $X$, and the adjacency matrix $A$ are available, with no access to any labels for ID or OOD nodes. This setting is consistent with the zero-shot OOD detection paradigm in the image domain.
@@ 3 Methodology
We address the zero-shot graph OOD detection problem by adapting GFM—a closed-world zero-shot classification method—to the open-set OOD setting. Notably, the original GFM model lacks mechanisms specifically designed for OOD detection. To enable OOD detection, we introduce all label names to the GFM and directly use its output logits for zero-shot OOD detection (§3.2). However, despite its effectiveness, the true OOD label space is often inaccessible in real-world settings. Therefore, we investigate zero-shot OOD detection with only ID labels as input (§3.3). This approach, however, suffers from limited expressiveness, as it relies solely on the absence of strong affinity with ID labels to infer OOD status—without any explicit OOD semantic reference, it struggles to detect semantically ambiguous or borderline cases. Therefore, we propose to prompt the LLM to generate a set of pseudo-OOD labels based on the initially unlabeled data, enabling the model to distinguish OOD samples in a more nuanced and fine-grained way. Our method detects OOD nodes by evaluating their affinities with both ID and pseudo-OOD labels (§3.4).
Figure 2: An overview of our framework, GLIP-OOD. Initially, the entire graph is unlabeled, and the labels of all nodes are unknown. GLIP-OOD leverages an LLM to generate semantically informative pseudo-OOD labels from unlabeled data, enabling the GFM to capture nuanced semantic boundaries between ID and OOD classes without requiring any labeled nodes.
@@ 3.1 Zero-shot Inference with Graph Foundation Model
Graph Encoder. For each node $v_{i}$, we construct its corresponding induced subgraph $G_{i}=(P_{i},X_{i},A_{i})$, and encode its information as follows:
$g_{i}=g_{\theta}(P_{i},X_{i},A_{i}),$ (1)
where $g_{\theta}$ denotes the Graph Transformer, such as GraphGPS [20], and $P_{i}$ represents the positional embeddings for the subgraph. Applying the mean pooling function yields the graph-level embedding $g_{i}\in\mathbb{R}^{1\times d}$ for the subgraph. In this way, we obtain an embedding $g_{i}$ for each input node $v_{i}$.
@@
where $h_{k}\in\mathbb{R}^{1\times d}$ is the embedding of each label sentence.
@@
$\hat{y}_{i}=\arg\max_{k}\mathbb{E}_{h_{k}}\,\text{sim}(g_{i},h_{k})$ (3)
@@ 3.2 Zero-shot Graph OOD Detection with All Labels
In this section, we provide all the label names to our framework for zero-shot OOD detection. Let $\mathcal{T}_{\text{ID}}=\{T_{1},T_{2},\dots,T_{K_{\text{ID}}}\}$ denote the set of ID label texts, and let $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$ denote the set of OOD label texts. Their corresponding embeddings are denoted as $\mathcal{H}_{\text{ID}}=\{h_{1},h_{2},\dots,h_{K_{\text{ID}}}\}$ and $\mathcal{H}_{\text{OOD}}=\{\bar{h}_{1},\bar{h}_{2},\dots,\bar{h}_{K_{\text{OOD}}}\}$, where $h_{k}=h_{\theta}(T_{k})$ and $\bar{h}_{k}=h_{\theta}(\bar{T}_{k})$ using the same text encoder $h_{\theta}$. We define the OOD score for node $v_{i}$ using its embedding $g_{i}$ and both ID and OOD label embeddings as follows:
@@
where $\text{sim}(g_{i},\mathcal{H})$ denotes the set of similarity scores between the node embedding $g_{i}$ and the embeddings in the label set $\mathcal{H}$, and $S^{*}$ is a fusion function that computes the OOD score based on these similarities. For example, we can first normalize all the similarity scores (including both ID and OOD labels), and then define the OOD score of node $v_{i}$ as one minus the sum of the normalized similarity scores between the node embedding $g_{i}$ and the sentence embeddings of the ID labels:
@@
where $K_{\text{ID}}$ is the number of ID classes. A higher OOD score indicates lower alignment with any known ID label.

### v2

# GLIP-OOD: Zero-Shot Graph OOD Detection with Graph Foundation Model

@@ Abstract
Out-of-distribution (OOD) detection is critical for ensuring the safety and reliability of machine learning systems, particularly in dynamic and open-world environments. In the vision and text domains, zero-shot OOD detection—which requires no training on in-distribution (ID) data—has advanced significantly through the use of large-scale pretrained models, such as vision-language models (VLMs) and large language models (LLMs). However, zero-shot OOD detection in graph-structured data remains largely unexplored, primarily due to the challenges posed by complex relational structures and the absence of powerful, large-scale pretrained models for graphs. In this work, we take the first step toward enabling zero-shot graph OOD detection by leveraging a graph foundation model (GFM). Our experiments show that, when provided only with class label names for both ID and OOD categories, the GFM can effectively perform OOD detection—often surpassing existing “supervised” OOD detection methods that rely on extensive labeled node data. We further address the practical scenario in which OOD label names are not available in real-world settings by introducing GLIP-OOD, a framework that uses LLMs to generate semantically informative pseudo-OOD labels from unlabeled data. These generated OOD labels allow the GFM to better separate ID and OOD classes, facilitating more precise OOD detection—all without any labeled nodes (only ID label names). To our knowledge, this is the first approach to achieve node-level graph OOD detection in a fully zero-shot setting, and it attains performance comparable to state-of-the-art supervised methods on four benchmark text-attributed graph datasets.
@@ 1 Introduction
Out-of-distribution (OOD) detection [33, 14, 9] is the task of identifying inputs that deviate significantly from the training distribution. As machine learning models are increasingly deployed in dynamic and safety-critical environments, the ability to reliably detect OOD instances has become important for maintaining robustness, improving trustworthiness, and preventing catastrophic failures. Applications span many domains, including medical diagnosis [25], autonomous driving [6], video analytics [11], and finance [2], where unseen or unexpected data is common and consequential.
@@
Graph OOD Detection and Current Landscape. Recently, OOD detection has gained traction in graph-structured data (known as graph OOD detection [29, 16, 31, 24]), where the goal is to detect anomalous or previously unseen nodes, edges, or subgraphs. It is important to the domains such as social networks [30], citation graphs, and e-commerce platforms, where new entities (e.g., emerging users, novel research topics, or innovative products) may appear during inference. Node-level graph OOD detection methods generally require a substantial number of labeled in-distribution (ID) nodes to train a graph neural network (GNN)-based classifier, which subsequently produces class probability distributions (logits) for all nodes in the graph [29, 16, 31]. OOD detection is then performed in a post-hoc manner by applying scoring functions [20]—such as energy-based methods [29, 14], maximum softmax probability [9], or uncertainty-based metrics [16]—to these logits. Nodes with high energy or entropy scores are considered likely to be OOD samples. However, these methods are fundamentally limited by their reliance on labeled ID data, which can be expensive and impractical to obtain in real-world graph scenarios. This raises a natural question: Can we perform graph OOD detection without relying on labeled nodes at all?
Figure 1: Radar plot (AUROC and AUPR) on four text-attributed graph datasets. GLIP-OOD-R uses real OOD class names, while GLIP-OOD-L employs LLM-generated pseudo-OOD names. With no node-level supervision, GLIP-OOD-R exceeds supervised OOD baselines, and GLIP-OOD-L remains competitive even without real OOD labels.
@@
Zero-Shot OOD Detection in Vision and Text. In computer vision and NLP areas, zero-shot OOD detection—which requires no training on ID data—has emerged as a compelling research direction. In the image domain, methods [26, 19, 5, 1, 15] leverage the capabilities of large-scale vision-language models (VLMs), e.g., CLIP [21], to detect OOD samples across a variety of ID datasets without requiring any labeled training examples. For example, CLIPN [26] enhances CLIP with the ability to distinguish between OOD and ID samples by leveraging both positive semantic prompts and negation-based prompts, enabling zero-shot OOD detection in the image domain. ZOC [5] extends CLIP by dynamically generating candidate unseen labels per test image and introducing a confidence score based on feature-space similarity to both seen and generated unseen labels. In the text domain, large language models (LLMs) have demonstrated impressive zero-shot capabilities by leveraging natural language prompts and contextual understanding to detect OOD content [34, 10]. The ability to perform effectively with limited data makes LLM-based OOD detection particularly attractive for real-world applications. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored due to the unique challenges posed by relational structures. More importantly, until now, there has been no widely acknowledged, powerful, large-scale foundation model—like VLMs and LLMs—for graphs.
Our Proposal: Zero-Shot OOD Detection in Graphs (GLIP-OOD). In this paper, we take the first step toward zero-shot graph OOD detection using a graph foundation model (GFM) [35, 12, 13, 3]. Similar to CLIP, the GFM does not inherently support OOD detection and instead matches any given node to one of the provided class labels. Below, we describe two scenarios: 1. GLIP-OOD-R (Ideal Scenario: All Class Names). When all class names (ID and OOD) are known, we supply only these label names to the GFM and derive the output logits to perform zero-shot OOD detection. Figure 1 shows that, surprisingly, with only these label names (GLIP-OOD-R), the GFM already surpasses many supervised methods that rely on numerous ID-labeled nodes. This result highlights the GFM’s strong intrinsic capacity for zero-shot OOD detection, given clearly defined ID and OOD classes.
2. GLIP-OOD-L (Practical Scenario: ID Names + Pseudo-OOD). More practically, the OOD classes are often unknown during training. We thus generate pseudo-OOD labels from unlabeled nodes to construct a pseudo-OOD label set alongside the known ID labels. Concretely, we sample a small subset of unlabeled nodes, ask an LLM to check if each node’s text matches any known ID category, and, if not, treat it as pseudo-OOD. The LLM then summarizes the node to produce a plausible OOD label name. This procedure yields new pseudo-OOD categories, which we feed back into the GFM for zero-shot OOD detection. Our approach requires no real OOD labels or node annotations, yet it can assign nodes to either known ID or newly synthesized OOD classes—enabling fine-grained OOD detection. As shown in Figure 1 (GLIP-OOD-L), the resulting performance is often comparable to fully supervised methods despite using no node-level supervision.
@@
  • Novel Setting. We are the first to study the problem of zero-shot graph OOD detection, providing a comprehensive setting and framework that detects OOD nodes without any node-level labels.
  • Zero-shot Solution. We propose GLIP-OOD, which leverages LLMs to generate pseudo-OOD labels, semantically distant from known ID labels but close to real OOD distributions in the transductive graph setting, then uses a GFM in an augmented label space for OOD detection.
  • Effectiveness. Across four benchmark text-attributed graph (TAG) datasets, our zero-shot method outperforms supervised approaches (when given all label names) and achieves performance comparable to those methods using only ID labels plus pseudo-OOD labels.
@@ 2 Preliminary
Our study focuses on OOD detection on TAGs. A TAG is represented as $G_{T}=(\mathcal{V},\mathbf{A},\mathcal{M},\mathbf{X})$, where the set of nodes is $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, and each node is associated with raw text. These text attributes =$\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$ are transformed into embeddings $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$ using a pretrained model [23]. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes the graph’s connectivity, where $\mathbf{A}[i,j]=1$ indicates an edge between nodes $v_{i}$ and $v_{j}$.
Node-level Graph OOD Detection. We have a labeled node set $\mathcal{V}_{I}$ and a large unlabeled node set $\mathcal{V}_{U}$. Each labeled node belongs to one of $K_{\text{ID}}$ known classes $\mathcal{Y}_{\text{ID}}=\{y_{k}\}_{k=1}^{K_{\text{ID}}}$, while an unlabeled node may belong to an unknown class not included in $\mathcal{Y}_{\text{ID}}$. The problem is framed in a semi-supervised, transductive setting, where we can access the full set of nodes during training but only a subset of class labels (ID classes). Our goal is to determine, for each node $v_{i}\in\mathcal{V}_{U}$, whether it is from the $K_{\text{ID}}$ ID classes or from the OOD unknown classes.
Zero-shot Graph OOD Detection with All Labels. We provide the model only with the ID and OOD class names and do not assume access to any node labels in the graph. That is, all nodes in the graph are unlabeled, and the model directly outputs an OOD score for each input node based solely on the label names, node attributes $\mathbf{X}$, and the adjacency matrix $\mathbf{A}$.
Zero-shot Graph OOD Detection. Only the names of ID classes, node attributes $\mathbf{X}$, and the adjacency matrix $\mathbf{A}$ are available, with no access to any labels for ID or OOD nodes. This setting is consistent with the zero-shot OOD detection paradigm in the image domain.
@@ 3 Methodology
Overview. We address the zero-shot graph OOD detection problem by adapting GFM—a closed-world zero-shot classification method—to the open-set OOD setting. Notably, the original GFM model lacks mechanisms specifically designed for OOD detection. To enable OOD detection, we introduce all label names to the GFM and directly use its output logits for zero-shot OOD detection (§3.2). However, despite its effectiveness, the true OOD label space is often inaccessible in real-world settings. Therefore, we investigate zero-shot OOD detection with only ID labels as input (§3.3). This approach, however, suffers from limited expressiveness: it infers OOD status solely from the absence of strong affinity with ID labels. Without any explicit semantic reference for OOD, the model struggles to identify semantically ambiguous or borderline cases. To address this, we propose prompting the LLM to generate a set of pseudo-OOD labels based on the initially unlabeled data, enabling the model to distinguish OOD samples in a more nuanced and fine-grained manner. Our method detects OOD nodes by evaluating their affinities with both ID and pseudo-OOD labels (§3.4).
Figure 2: GLIP-OOD overview. Initially, the entire graph is unlabeled, and the labels of all nodes are unknown. GLIP-OOD leverages an LLM to generate semantically informative pseudo-OOD labels from unlabeled data, enabling the GFM to capture nuanced semantic boundaries between ID and OOD classes without requiring any labeled nodes.
@@ 3.1 Zero-shot Inference with Graph Foundation Model
Graph Encoder. For each node $v_{i}$, we construct its corresponding induced subgraph $G_{i}=(\mathbf{P}_{i},\mathbf{X}_{i},\mathbf{A}_{i})$, where $\mathbf{P}_{i}$ denotes the positional encodings, $\mathbf{X}_{i}$ is the feature matrix of the subgraph (e.g., the $k$-hop neighborhood centered at node $v_{i}$), and $\mathbf{A}_{i}$ is the adjacency matrix of the subgraph. The subgraph is then encoded using a Graph Transformer $g_{\theta}$, such as GraphGPS [22], as follows:
$g_{i}=g_{\theta}(\mathbf{P}_{i},\mathbf{X}_{i},\mathbf{A}_{i}),$ (1)
where $g_{i}\in\mathbb{R}^{1\times d}$ denotes the graph-level embedding obtained via mean pooling over node representations. This results in a fixed-dimensional embedding $g_{i}$ for each input node $v_{i}$.
@@
where $h_{\phi}$ denotes the MiniLM encoder and $h_{k}\in\mathbb{R}^{1\times d}$ is the embedding of each label sentence.
@@
$\hat{y}_{i}\;=\;\underset{k\in\{1,\dots,K\}}{\operatorname*{arg\,max}}\operatorname{sim}\!\bigl{(}g_{i},h_{k}\bigr{)},$ (3)
@@ 3.2 Zero-shot Graph OOD Detection with All Labels
In this section, we provide all the label names to our framework for zero-shot OOD detection. Let $\mathcal{T}_{\text{ID}}=\{T_{1},T_{2},\dots,T_{K_{\text{ID}}}\}$ denote the set of ID label texts, and let $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$ denote the set of OOD label texts. Their corresponding embeddings are denoted as $\mathcal{H}_{\text{ID}}=\{h_{1},h_{2},\dots,h_{K_{\text{ID}}}\}$ and $\mathcal{H}_{\text{OOD}}=\{\bar{h}_{1},\bar{h}_{2},\dots,\bar{h}_{K_{\text{OOD}}}\}$, where each embedding is computed using the same text encoder $h_{\phi}$ defined in Eqn. 2, with $h_{k}=h_{\phi}(T_{k})$ and $\bar{h}_{k}=h_{\phi}(\bar{T}_{k})$. We define the OOD score for node $v_{i}$ using its embedding $g_{i}$ and both ID and OOD label embeddings as follows:
@@
where $\text{sim}(g_{i},\mathcal{H})$ denotes the set of similarity scores between the node embedding $g_{i}$ and the embeddings in the label set $\mathcal{H}$, and $S^{*}$ is a fusion function that computes the OOD score based on these similarities. For example, we can normalize all similarity scores (from both ID and OOD labels), and define the OOD score of node $v_{i}$ as one minus the sum of the normalized similarity scores between $g_{i}$ and the ID label embeddings:
@@
where $K_{\text{ID}}$ is the number of ID classes, and we use the cosine similarity function in this paper. A higher OOD score indicates lower alignment with any known ID label.